OpenStudy (anonymous):

What is the minimum or maximum? y= -x^2-x-1

OpenStudy (jdoe0001):

is a parabola, the leading coefficient, that is from \(\bf -x^2\) is negative, so it's going downwards so it'll look like  the HIGHEST point will be where the vertex is at you can get the y-coordinate for the vertex with \(\bf c-\cfrac{b^2}{4a}\)

OpenStudy (anonymous):

i already got that. but i cant figure out the equation for y coordinate

OpenStudy (jdoe0001):

sorry was... a bit elsewhere.... \(\bf \begin{array}{llll} &a&b&c\\ y= -x^2-x-1 \implies y= &-1x^2&-1x&-1 \end{array}\\ c-\cfrac{b^2}{4a}\)

OpenStudy (jdoe0001):

a, b and c, are just the coefficients

OpenStudy (anonymous):

i got .25?

OpenStudy (jdoe0001):

well, \(\bf c-\cfrac{b^2}{4a} \implies (-1)-\left(\cfrac{(-1)^2}{4(-1)}\right) \implies (-1)-\left(\cfrac{1}{-4}\right) \\\quad \\ \implies -1 + \cfrac{1}{4} \implies -\cfrac{4}{4}+\cfrac{1}{4} \implies -\cfrac{3}{4} = 0.75\)

OpenStudy (jdoe0001):

woops, -0.75 for that matter, missed the - there

OpenStudy (jdoe0001):

\(\bf -\cfrac{3}{4} = -0.75\)